Let A= {1, 2, 3, 4}, and let R And S be the relations on A As follow: R = {(1,1),(1, 4),(3,2),(4,2),(4,3)} S = { (1,1),(1,2),(2,2),(3,3),(4,2),(4,4)) Then find out: 1. M𝑅𝑐 2. 𝑀𝑠̅ 3. M (𝑅βˆͺ S)


1
Expert's answer
2021-11-02T15:36:25-0400

1.

"R^c=A\times A-R"

"M(R^c)=\begin{pmatrix}
 0 & 1&1&0 \\
 1 & 1&1&1\\
1 & 0&1&1\\
1 & 0&0&1
\end{pmatrix}"


2.

"\overline{S}=A\times A-S"

"M(\overline{S})=\begin{pmatrix}
 0 & 0&1&1 \\
 1 & 0&1&1\\
1 & 1&0&1\\
1 & 0&1&0
\end{pmatrix}"


3.

"R\cup S=\{ (1,1),(1,2),(1,4),(2,2),(3,2),(3,3),(4,2),(4,3),(4,4)\}"

"M(R\cup S)=\begin{pmatrix}
 1 & 1&0&1 \\
 0 & 1&0&0\\
0 & 1&1&0\\
0 & 1&1&1
\end{pmatrix}"